Holding constant the cost of the basket in each country, compute the real exchange rates that would result from the two nominal exchange rates in the following table. Cost of Basket in U.S. (Dollars) 18 18 Cost of Basket in Japan (Yen) 570 570 Nominal Exchange Rate (Yen per dollar) 95.00 76.00 Real Exchange Rate (Baskets of Japanese goods per basket of U.S. goods)
Answer #1

Solution-

Cost of Basket in Cost of basket in Nominal Exchange Real Exchange

U.S. Japan rate rate

18 570 95.00 3

18 570 76.00 2.4

Reason-

18 x 95 = 1710

Real exchange rate = 1710 / 570 = 3

18 x 76 1368

Real exchange rate = 1368 / 570 = 2.4